# Finite groups and arc-transitive maps of square-free Euler characteristic

### v2

@@ 3. Examples of square-free Euler characteristic regular maps
Thus the map ${\mathcal{M}}_{n}$ has valency $4$ and face length $2n$. Let ${\it\Gamma}={\sf Cos}(G,H,HzH)$ be the underlying graph of ${\mathcal{M}}_{n}$. Then $G=\langle a,b\rangle{:}H$, and so $\langle a,b\rangle$ is regular on the vertex set of $\it\Gamma$}. Hence $\it\Gamma$} is a (connected and undirected) Cayley graph ${\sf Cay}(K,S)$ of valency 4, where $K=\langle a,b\rangle={\sf Z}_{n}^{2}$. It follows that $S=\{x,x^{-1},y,y^{-1}\}$ with $\langle x,y\rangle=\langle a,b\rangle$, and so ${\it\Gamma}={\bf C}_{n}\square{\bf C}_{n}$.
The Euler characteristic
$\chi({\mathcal{M}}_{n})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=8n^{2}\left({1\over 8}-{1\over 4}+{1\over 4n}\right)={-n(n-2).}$
According to [8], there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. Further, by a theorem of Chen [3], there indeed exist infinitely many primes $p$ such that p2p-2 is square-free, then so is p(p2)-p(p-2). $\Box$
Here is some data such that p(p2)-p(p-2) is square-free, with $p$ prime.
$\begin{array}[]{c|cccccccc}\hline\cr p&3&5&7&13&17&19&23&\cdots\\ -p(p-2)&-3&-3.5&-5.7&-11.13&-3.5.17&-17.19&-3.7.23$$\cdots\\ \hline\cr\end{array}$
@@ 4. $p$-Groups
In this section, we study $p$-groups satisfying Hypothesis 2.2, i.e., $p$-groups with a cyclic or dihedral subgroup of index $p$. Note that the case for $p=2$ is dealt with in the previous work [10], which is introduced below.
@@
([10], Thm. 1.1) Let $G$ be a $2$-group which has a cyclic or dihedral subgroup of index $2$. Then $G$ is one of the following groups. Further, either ${\sf Aut}(G)$ is a $2$-group, or $(G,{\sf Aut}(G))$ is one of $({\sf Z}_{2}^{2},{\rm S}_{3}),\,({\sf Z}_{2}^{3},{\rm GL}_{3}(2)),\,({\rm Q}_{8},{\rm S}_{4}),\,({\rm Q}_{8}\circ{\sf Z}_{4},{\rm S}_{4}\times{\sf Z}_{2})$.
@@
    If $G$ has a reversing triple, then $G={\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$, ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$ or ${\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4}$, where $\ell\geqslant 1;$ further, if $G$ has a regular triple, then $G\neq{\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4};$
@@
For the odd primes $p$, we have the following characterization.
@@
    $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ is abelian, and
@@
      if $\ell>1$, then ${\sf Aut}(G)$ is solvable, and letting $G=\langle a,b\rangle$, then any Hall $p^{\prime}$-subgroup of ${\sf Aut}(G)$ is conjugate to $({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1};$
@@
      if $\ell=1$, then ${\sf Aut}(G)={\rm GL}(2,p)$, and further, any subgroup $M\leqslant{\rm GL}(2,p)$ of order coprime to $2p$ is an abelian group.
@@
    $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ is non-abelian with $a^{b}=a^{p^{\ell-1}+1}$, $\ell\geqslant 2$, and ${\sf Aut}(G)$ is solvable such that any Hall $p^{\prime}$-subgroups of ${\sf Aut}(G)$ is conjugate to $\langle\rho\rangle={\sf Z}_{p-1}$, where $a^{\rho}=a^{\xi}\not=a$ with $\xi\,\,\big|\,\,p-1$, and $b^{\rho}=b$.
@@
(ii) Suppose $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ with $\ell>1$. Then by [1, Thm. 3.2], ${\sf Aut}(G)$ is of order
$\begin{array}[]{ccl}|{\sf Aut}(G)|&=&|{\sf Aut}({\sf Z}_{p^{\ell}})|\cdot|{\sf Aut}({\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p^{\ell}},{\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p},{\sf Z}_{p^{\ell}})|\\ &=&(p-1)p^{\ell-1}\cdot(p-1)\cdot p\cdot p=p^{\ell+1}(p-1)^{2}.\end{array}$
More explicitly, let $G=\langle a\rangle\times\langle b\rangle$ with $|a|=p^{\ell}$, $|b|=p$. Then $\Phi(G)=\langle a^{p}\rangle$. Let $\overline{G}=G/\Phi(G)$, and consider the homomorphism
@@
exactly consists of those elements of $G$ of order less than $p^{\ell}$. Moreover, it follows from ${\sf Aut}(G)/{\rm Ker}(\sigma)\cong{\rm Im}(\sigma)$ that $|{\rm Im}(\sigma)|=p(p-1)^{2}=|P|$, and so ${\rm Im}(\sigma)=P$.
Now both ${\rm Ker}(\sigma)$ and ${\rm Im}(\sigma)$ are solvable, so is ${\sf Aut}(G)$. Note that the following subgroup
$X=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1}$
is a Hall $p^{\prime}$-subgroup of $G$. Then any other Hall $p^{\prime}$-subgroup is conjugate to $X$.
@@
(iii) Suppose $G={\sf Z}_{p}\times{\sf Z}_{p}$. Then ${\sf Aut}(G)={\rm GL}(2,p)$ is of order $p(p+1)(p-1)^{2}$, with prime divisors less than or equal to $p$. Further, let $M\leqslant{\rm GL}(2,p)$ be a subgroup of order coprime to $2p$, and let $N=M/{\bf Z}(M)$. Then
$N=M/{\bf Z}(M)\leqslant M/({\bf Z}({\rm GL}(2,p))\cap M)\lesssim{\rm PGL}(2,p).$
Further, since ${\rm PSL}(2,p)N\leqslant{\rm PGL}(2,p)$, we have
$|N/({\rm PSL}(2,p)\cap N)|\leqslant|{\rm PGL}(2,p)|/|{\rm PSL}(2,p)|=2.$
@@
It then follows by $2\nmid|N|$ that $N$ is indeed embedded into ${\rm PSL}(2,p)$. Note that the subgroups of ${\rm PSL}(2,p)$ are all explicitly known, by check which a subgroup of order coprime to $2p$ must be a cyclic group. Then $M$ is an abelian group, since $N=M/{\bf Z}(M)$ is cyclic.
@@
(iv) Suppose that $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ with $a^{b}=a^{p^{\ell-1}+1}$ is non-abelian, $\ell\geqslant 2$. The automorphism group ${\sf Aut}(G)$ is described in [2, Thm. 3.1 & 3.2], by which we have: $|{\sf Aut}(G)|=(p-1)p^{\ell+1}$, ${\sf Aut}(G)$ is solvable with the Sylow $p$-subgroup normal, and $X=\langle\rho\rangle={\sf Z}_{p-1}$ is a Hall $p^{\prime}$-subgroup of $G$, where $\rho$ stabilizes $\langle a\rangle$ and centralizes $\langle b\rangle$, that is, $a^{\rho}=a^{\xi}\not=a$ with $\xi\,\,\big|\,\,p-1$, and $b^{\rho}=b$. $\Box$
@@
In this section, we study the way how the $p$-groups described in Section 4 compose a larger solvable group, and give a proof of Theorem 1.1
Let $G$ be a finite group, with $\{p_{1},\ldots,p_{s}\}$ the prime divisors of $|G|$. For $p=p_{i}$, denote by ${\bf O}_{p}(G)$ the maximal normal $p$-subgroup of $G$, or equivalently, the intersection of all Sylow $p$-subgroups of $G$. Let $F={\rm Fit}(G)$ be the Fitting subgroup of $G$, which is the largest nilpotent normal subgroup. Then
@@
Note that both ${\bf O}_{p}(G)$ and $F$ are characteristic in $G$. Further, if $G$ is solvable, then ${\bf C}_{G}(F)\leqslant F$, and so
@@
Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2. Let $p$ be the largest prime divisor of $|G|$. Then a Sylow $p$-subgroup $G_{p}$ is normal in $G$, and so $G=G_{p}{:}G_{p^{\prime}}$.
@@
Let $\overline{G}=G/{\bf O}_{p}(G)$, and let $F={\rm Fit}(\overline{G})$. Notice that ${\bf O}_{p}(\overline{G})=1$. We have $F=F_{p_{1}}\times\cdots\times F_{p_{s}}$ with all $p_{i}<p$. Now,
$\overline{G}/F\lesssim{\sf Out}(F)={\sf Out}(F_{p_{1}})\times\cdots\times{\sf Out}(F_{p_{s}}),$
and so $p\nmid|\overline{G}/F|$ by Lemma 4.3. Thus, $p\nmid|\overline{G}|$, so that, $G_{p}={\bf O}_{p}(G)\lhd G$. $\Box$
@@
Let $G$ be a $\{p,q\}$-group satisfying Hypotheses 2.2, where $p>q$ are both odd primes, and $q\nmid p^{2}-1$. Then $G=G_{p}\times G_{q}$.
@@
By Lemma 5.1, $G=G_{p}{:}G_{q}$. Since $G$ is solvable, we obtain that $G_{q}/{\bf C}_{G_{q}}(G_{p})\cong G/(G_{p}{\bf C}_{G_{q}}(G_{p}))=G/(G_{p}{\bf C}_{G}(G_{p}))\leq{\sf Out}(G_{p})$. By Lemma 4.3, any prime divisor of $|{\sf Out}(G_{p})|$ divides $p(p^{2}-1)$. However, $q\nmid p(p^{2}-1)$ by our assumption, and we conclude that $G_{q}/{\bf C}_{G_{q}}(G_{p})=1$, so that $G_{q}={\bf C}_{G_{q}}(G_{p})$ and $G=G_{p}\times G_{q}$. $\Box$
Before giving a general characterization of solvable groups of odd order, we need the following technical lemma.
@@
Let $H={\sf Z}_{p^{\ell}}$ be a cyclic group and $x$ be an automorphism of $H$ with order coprime to $p$. Let $X=H{:}\langle x\rangle$. Then, (i) $H\leqslant X^{\prime};$ (ii) $H\cap{\bf Z}(X)=1$.
@@
By Lemma 4.3 (1), the order of $x$ divides p1p-1. Let $H=\langle a\rangle$ and suppose
If $p\mid m-1$, then by LTE (Lift the Exponent Lemma) we have $p^{\ell}\mid m-1$, a contradiction. Hence $p\nmid m-1$. Thus, (i) the commutator $a,x]=a^{-1}a^{x}=a^{m-1}$ can generate $H$, so that, $H\leqslant X^{\prime}$; (ii) for any $1\leqslant k<p^{\ell}$, $(a^{k})^{x}=a^{km}\neq a^{k}$, so that, $H\cap{\bf Z}(X)=1$. $\Box$
Lemma 5.4.
Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2. Then G=C:DG=C{:}D, where $C$ is abelian, $D$ is nilpotent, and $C\cap{\bf Z}(G)=1$.
Proof.
Let $n$ be the number of prime divisors of $|G|$. The proof proceeds by induction on $n$. For $n=1$, the lemma holds with $C=1$, $D=G$. Suppose $n\geqslant 2$, and assume inductively that the lemma holds for groups of order with less prime divisors.
@@
with $A$ abelian, $B$ nilpotent, and $A\cap{\bf Z}(A{:}B)=1$, so that, $B$ does not centralize any Sylow subgroup of $A$. Further, if $G=G_{p}\times G_{p^{\prime}}$, then
$G=G_{p}{\times}(A{:}B)=(G_{p}{\times}A){:}B=A{:}(G_{p}\times B),$

### v3

@@ 3. Examples of Square-free Euler Characteristic Regular Maps
Thus, $(|V|,|E|,|F|)=(\frac{|G|}{|H|},\frac{|G|}{4},\frac{|G|}{|L|})=(n^{2},2n^{2},2n)$, and the Euler characteristic
$\chi({\mathcal{M}}_{n})=n^{2}-2n^{2}+2n=-n(n-2).$
It is similar to Lemma 3.2 to prove that, there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. Further, by a theorem of Chen [3], there indeed exist infinitely many primes $p$ such that p2p-2 is square-free, then so is p(p2)-p(p-2). Here are some examples :
$\begin{array}[]{|c|c|c|c|c|c|c|c|c|}\hline\cr p&3&5&7&13&17&19&23&\cdots\\ \hline\cr-p(p-2)&-3&-3.5&-5.7&-11.13&-3.5.17&-17.19&-3.7.23$$\cdots\\ \hline\cr\end{array}$
At last, the valency of ${\mathcal{M}}_{n}$ is $\frac{2|E|}{|V|}=4$, and the face length is $\frac{2|E|}{|F|}=2n$. Let ${\it\Gamma}={\sf Cos}(G,H,HJ)$ be the underlying graph of ${\mathcal{M}}_{n}$, and let $K=\langle a,b\rangle={\sf Z}_{n}^{2}\leqslant G$. Note that, G=K:HG=K{:}H, so $K$ is regular on the vertex set of $\it\Gamma$}, and $\it\Gamma$} is thus a connected Cayley graph ${\sf Cay}(K,S)$ of valency 4. It follows that $S=\{x,x^{-1},y,y^{-1}\}$ with $\langle x,y\rangle=\langle a,b\rangle$, and ${\it\Gamma}={\sf Cay}(K,S)\cong{\bf C}_{n}\square{\bf C}_{n}$. $\Box$
@@ 4. $p$-Groups
In this section, we study $p$-groups satisfying Hypothesis 2.2, i.e., $p$-groups with a cyclic or dihedral subgroup of index $p$. Note that, the case for $p=2$ is dealt with in the previous work [10], which is introduced below.
@@
([10], Thm. 1.1) Let $G$ be a $2$-group which has a cyclic or dihedral subgroup of index $2$. Then one of the following holds. Further, either ${\sf Aut}(G)$ is a $2$-group, or $(G,{\sf Aut}(G))$ is one of $({\sf Z}_{2}^{2},{\rm S}_{3}),\,({\sf Z}_{2}^{3},{\rm GL}_{3}(2)),\,({\rm Q}_{8},{\rm S}_{4}),\,({\rm Q}_{8}\circ{\sf Z}_{4},{\rm S}_{4}\times{\sf Z}_{2})$.
@@
    If $G$ has a reversing triple, then $G={\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$, ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$ or ${\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4}$, where $\ell\geqslant 1;$ further, if $G$ has a regular triple, then $G\neq{\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4}.$
@@
The case for $p>2$ is dealt with in the following lemma.
@@
    $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}=\langle a,b\rangle$ is abelian, and
@@
      if $\ell>1$, then ${\sf Aut}(G)$ is solvable of order $p^{\ell+1}(p-1)^{2}$, with an abelian Hall $p^{\prime}$-subgroup ${\sf Aut}(G)_{p^{\prime}}=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1};$
@@
      if $\ell=1$, then ${\sf Aut}(G)={\rm GL}(2,p)$ is of order $p(p+1)(p-1)^{2}$, with any subgroup $M\leqslant{\rm GL}(2,p)$ of order coprime to $2p$ being abelian.
@@
    $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ is non-abelian with $a^{b}=a^{p^{\ell-1}+1}$, $\ell\geqslant 2$, and ${\sf Aut}(G)$ is solvable of order $(p-1)p^{\ell+1}$, with a cyclic Hall $p^{\prime}$-subgroup ${\sf Aut}(G)_{p^{\prime}}=\langle\rho\rangle={\sf Z}_{p-1}$, where $a^{\rho}=a^{\xi}\not=a$ and $b^{\rho}=b$.
@@
(ii) Suppose $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}=\langle a,b\rangle$, with $|a|=p^{\ell}>p$, $|b|=p$. By [1, Thm. 3.2],
$\begin{array}[]{ccl}|{\sf Aut}(G)|&=&|{\sf Aut}({\sf Z}_{p^{\ell}})|\cdot|{\sf Aut}({\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p^{\ell}},{\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p},{\sf Z}_{p^{\ell}})|\\ &=&(p-1)p^{\ell-1}\cdot(p-1)\cdot p\cdot p\\ &=&p^{\ell+1}(p-1)^{2}.\end{array}$
In particular, ${\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle)\leqslant{\sf Aut}(G)$, so the subgroup
$({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}\cong({\sf Z}_{(p-1)p^{\ell-1}}\times{\sf Z}_{p-1})_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1}$
is an abelian Hall $p^{\prime}$-subgroup of ${\sf Aut}(G)$ of order $(p-1)^{2}$.
Let $\Phi(G)=\langle a^{p}\rangle$ be the Frattini subgroup of $G$, and let $\overline{G}=G/\Phi(G)$. Consider the homomorphism
@@
exactly consists of those elements of $G$ of order less than $p^{\ell}$. Moreover, ${\rm Im}(\sigma)=P$, since $|{\rm Im}(\sigma)|=|{\sf Aut}(G)|/|{\rm Ker}(\sigma)|=p(p-1)^{2}=|P|$.
Now, both ${\rm Ker}(\sigma)$ and ${\rm Im}(\sigma)$ are solvable, so is ${\sf Aut}(G)$.
@@
(iii) Suppose $G={\sf Z}_{p}\times{\sf Z}_{p}$. Then, ${\sf Aut}(G)={\rm GL}(2,p)$ is of order $(p^{2}-1)(p^{2}-p)=p(p+1)(p-1)^{2}$, with prime divisors less than or equal to $p$. Let $M\leqslant{\rm GL}(2,p)$ be a subgroup of order coprime to $2p$, and let $N=M/{\bf Z}(M)$. Then,
$N=M/{\bf Z}(M)\leqslant M/({\bf Z}({\rm GL}(2,p))\cap M)\lesssim{\rm PGL}(2,p),$
@@
and, in fact, $N$ is embedded into ${\rm PSL}(2,p)$, because $2\nmid|N|$, and meanwhile, ${\rm PSL}(2,p)$ is a index 2 normal subgroup of ${\rm PGL}(2,p)$. The subgroups of ${\rm PSL}(2,p)$ are all explicitly known, by check which a subgroup of order coprime to $2p$ must be a cyclic group. Thus, $N=M/{\bf Z}(M)$ is cyclic, so that, $M$ is abelian.
@@
(iv) Suppose $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ with $a^{b}=a^{p^{\ell-1}+1}$, $\ell\geqslant 2$. By [2, Thm. 3.1], $|{\sf Aut}(G)|=(p-1)p^{\ell+1}$, and hence the Sylow subgroup ${\sf Aut}(G)_{p}$ is normal in ${\sf Aut}(G)$. By [2, Thm. 3.2], there exists a cyclic Hall $p^{\prime}$-subgroup ${\sf Aut}(G)_{p^{\prime}}=\langle\rho\rangle={\sf Z}_{p-1}$, where $\rho$ stabilizes $\langle a\rangle$ and centralizes $\langle b\rangle$ as $a^{\rho}=a^{\xi}\not=a$ and $b^{\rho}=b$. Then, ${\sf Aut}(G)={\sf Aut}(G)_{p}{:}{\sf Aut}(G)_{p^{\prime}}$ is solvable. $\Box$
The following corollary will be used in the next Section.
Corollary 4.4.
Let $H$ be a $3$-group satisfying Hypothesis 2.2. Then the following statements hold.
  • (1)
    Let $\sigma\in{\sf Aut}(H)$ be an involution. Then $H{:}\langle\sigma\rangle$ is isomorphic to one of ${\rm D}_{2.3^{\ell}}$, ${\rm D}_{2.3^{\ell}}\times{\sf Z}_{3}$, ${\sf Z}_{3^{\ell}}\times{\rm D}_{6}$, $({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}$ and ${\sf Z}_{3^{\ell}}{:}{\sf Z}_{6}$.
  • (2)
    If ${\sf Aut}(H)$ contains a subgroup $X$ such that $X\cong{\sf Z}_{2}^{2}$, then $H={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$, and $H{:}X={\rm D}_{2.3^{\ell}}\times{\rm D}_{6}$.
  • (3)
    The group ${\sf Out}(H)$ contains no subgroups isomorphic to ${\rm A}_{4}$.
Proof.
By Lemma 4.3, $H$ is one of ${\sf Z}_{3^{\ell}}$, ${\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$ and ${\sf Z}_{3^{\ell}}{:}{\sf Z}_{3}$. Let $\sigma\in{\sf Aut}(H)$ be an involution.
(i) If $H={\sf Z}_{3^{\ell}}$, then ${\sf Aut}(H)={\sf Z}_{2.3^{\ell-1}}$ contains only one involution, $H{:}\langle\sigma\rangle={\rm D}_{2.3^{\ell}}$, and ${\sf Out}(H)={\sf Aut}(H)$ contains no subgroups isomorphic to ${\sf Z}_{2}^{2}$ or ${\rm A}_{4}$.
(ii) If $H={\sf Z}_{3^{\ell}}{:}{\sf Z}_{3}=\langle a\rangle{:}\langle b\rangle$ is non-abelian as given in Lemma 4.3 (3), then ${\sf Aut}(H)$ is solvable of order $2.3^{\ell+1}$, so $\langle\sigma\rangle$ is a Hall $3^{\prime}$-subgroup of ${\sf Aut}(H)$ of order 2, and it is conjugate to the following one:
$\langle\rho\rangle={\sf Z}_{2}$, with $(a,b)^{\rho}=(a^{-1},b)$.
Thus, $H{:}\langle\sigma\rangle\cong H{:}\langle\rho\rangle={\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3}={\sf Z}_{3^{\ell}}{:}{\sf Z}_{6}$. Meanwhile, since $|{\sf Aut}(H)|$ is not divisible by 4, ${\sf Aut}(H)$ has no subgroups isomorphic to ${\sf Z}_{2}^{2}$, and ${\sf Out}(H)$ has no subgroups isomorphic to ${\rm A}_{4}$.
(iii) If $H={\sf Z}_{3}\times{\sf Z}_{3}$, then ${\sf Aut}(H)={\sf Out}(H)$, and ${\sf Aut}(H)={\rm GL}_{2}(3)\cong{\rm Q}_{8}{:}{\rm S}_{3}$ is solvable, containing one central involution and twelve non-central involutions, with the non-central ones conjugate to the element with diagonal matrix (1,1)(-1,1). Hence, $H{:}\langle\sigma\rangle$ is isomorphic to $({\sf Z}_{3}\times{\sf Z}_{3}){:}{\sf Z}_{2}$ or ${\rm D}_{6}\times{\sf Z}_{3}$, where in the former case, ${\bf Z}({\rm GL}_{2}(3))={\sf Z}_{2}$ acts on ${\sf Z}_{3}\times{\sf Z}_{3}$ by mapping each element to its inverse.
Let $X\leqslant{\sf Aut}(H)$ be a subgroup with $X\cong{\sf Z}_{2}^{2}$. Then, $X\cap{\rm Q}_{8}>1$, so $X$ contains the central involution with diagonal matrix (1,1)(-1,-1), and $X$ contains other two non-central involutions conjugate to (1,1)(-1,1) and (1,1)(1,-1). Hence $H{:}X={\rm D}_{6}\times{\rm D}_{6}$. Moreover, if ${\sf Aut}(H)$ contains a subgroup isomorphic to ${\rm A}_{4}\cong{\sf Z}_{2}^{2}{:}{\sf Z}_{3}$, then ${\sf Z}_{3}$ acts on ${\sf Z}_{2}^{2}$ by moving the central involution, not possible.
(iv) At last, suppose $H={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}=\langle a,b\rangle$, where $|a|=3^{\ell}>3$, $|b|=3$. Then, ${\sf Aut}(H)={\sf Out}(H)$. Let $X\leqslant{\sf Aut}(H)$ be a subgroup with $X\cong{\sf Z}_{2}^{2}$. By Lemma 4.3 (2), ${\sf Aut}(H)$ is solvable of order $4.3^{\ell+1}$, so $X$ is a Hall $3^{\prime}$-subgroup of ${\sf Aut}(H)$ of order $4$, and it is conjugate to the following one:
$M=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{3^{\prime}}={\sf Z}_{2}\times{\sf Z}_{2}=\langle x,y\rangle,$
where $(a,b)^{x}=(a^{-1},b)$ and $(a,b)^{y}=(a,b^{-1})$. Thus, $H{:}X\cong H{:}M={\rm D}_{2.3^{\ell}}\times{\rm D}_{6}$. In particular, the involution $\sigma\in{\sf Aut}(H)$ is contained in some Hall $3^{\prime}$-subgroup, thus conjugate to $x$, $y$ or $xy$; respectively, $H{:}\langle\sigma\rangle$ is isomorphic to ${\rm D}_{2.3^{\ell}}\times{\sf Z}_{2}$, ${\sf Z}_{3^{\ell}}\times{\rm D}_{6}$ or $({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}$.
At last, let $g\in{\bf N}_{{\sf Aut}(H)}(M)$. Note that, $M$ stabilizes subgroups $\langle a\rangle$ and $\langle b\rangle$, and meanwhile it cannot stabilize
Now, $M=M^{g}$ stabilizes $\langle a\rangle^{g}$ and $\langle b\rangle^{g}$, which implies $\langle a\rangle^{g}=\langle a\rangle$, $\langle b\rangle^{g}=\langle b\rangle$, so that, ${\bf N}_{{\sf Aut}(H)}(M)\leqslant{\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle)$ is abelian. Since any subgroup $Y\leqslant{\sf Aut}(H)$ with $Y\cong{\sf Z}_{2}^{2}$ is conjugate to $M$, the normalizer ${\bf N}_{{\sf Aut}(H)}(Y)$ is conjugate to ${\bf N}_{{\sf Aut}(H)}(M)$, so it is abelian, too. Consequently, ${\sf Aut}(H)$ has no subgroup isomorphic to ${\rm A}_{4}$. $\Box$
@@
In this section, we give a proof of Theorem 1.1, by studying the way how the $p$-groups compose a larger solvable group.
Let $G$ be a finite group, with $\{p_{1},\ldots,p_{s}\}$ the prime divisors of $|G|$. For $p=p_{i}$, denote by ${\bf O}_{p}(G)$ the maximal normal $p$-subgroup of $G$, or equivalently, the intersection of all Sylow $p$-subgroups of $G$. Let $F={\rm Fit}(G)$ be the Fitting subgroup of $G$, which is the unique largest nilpotent normal subgroup. Then,
@@
Note that, both ${\bf O}_{p}(G)$ and ${\rm Fit}(G)$ are characteristic in $G$. Further, if $G$ is solvable, then ${\bf C}_{G}(F)\leqslant F$, and so
@@
Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2. Let $p$ be the largest prime divisor of $|G|$. Then:
  • (1)
    a Sylow p-subgroup $G_{p}$ is normal in $G$, and hence $G=G_{p}{:}G_{p^{\prime}}\,;$
  • (2)
    if $G$ is a $\{p,q\}$-group, with q a prime such that $q<p$ and $q\nmid p^{2}-1$, then $G=G_{p}\times G_{q}$.
@@
(1) Let $\overline{G}=G/{\bf O}_{p}(G)$. Note that, $\overline{G}$ has no normal $p$-subgroup. Thus, $F:={\rm Fit}(\overline{G})=F_{p_{1}}\times\cdots\times F_{p_{s}}$ with all $p_{i}<p$, and
$\overline{G}/F\lesssim{\sf Out}(F)={\sf Out}(F_{p_{1}})\times\cdots\times{\sf Out}(F_{p_{s}}).$
By Lemma 4.3, $p\nmid|{\sf Out}(F_{p_{i}})|$ as $p>p_{i}$. Then, $p\nmid|\overline{G}/F|$. It follows that $p\nmid|\overline{G}|$, so $G_{p}={\bf O}_{p}(G)\lhd G$ and $G=G_{p}{:}G_{p^{\prime}}$.
(2) Suppose that $G$ is a $\{p,q\}$}-group with $p>q$ and $q\nmid p^{2}-1$. By (1), we have $G=G_{p}{:}G_{q}$. Meanwhile, $H:=G/{\bf O}_{q}(G)$ has no normal $q$-subgroup, so ${\rm Fit}(H)$ is a $p$-group, and $H/{\rm Fit}(H)\lesssim{\sf Out}({\rm Fit}(H))$. By Lemma 4.3, any prime divisor of $|{\sf Out}({\rm Fit}(H))|$ is a divisor of $p(p^{2}-1)$. Then, $q\nmid|H/{\rm Fit}(H)|$ as $q\nmid p(p^{2}-1)$. It follows that $q\nmid|H|$, so $G_{q}={\bf O}_{q}(G)\lhd G$ and $G=G_{p}\times G_{q}$. $\Box$
Before giving a general characterization of solvable groups of odd order, we need the following fact.
@@
Let $H={\sf Z}_{p^{\ell}}$, $1\neq x\in{\sf Aut}(H)$ be an automorphism with $\gcd(|x|,p)=1$, and $X=H{:}\langle x\rangle$. Then, (i) $H\leqslant X^{\prime}\,;$ (ii) $H\cap{\bf Z}(X)=1$.
@@
Let $H=\langle a\rangle$. By Lemma 4.3 (1), $|x|$ divides p1p-1, and thus we can suppose
If $p\mid m-1$, then by LTE (Lift the Exponent Lemma), we have $p^{\ell}\mid m-1$, so $a^{x}=a^{m}=a$ and $x=1$, a contradiction. Hence $p\nmid m-1$. Thus, (i) the commutator $a,x]=a^{-1}a^{x}=a^{m-1}$ can generate $H$, so that, $H\leqslant X^{\prime}$ ; (ii) for any $1\leqslant k<p^{\ell}$, $(a^{k})^{x}=a^{km}\neq a^{k}$, so that, $H\cap{\bf Z}(X)=1$. $\Box$
@@
Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2. Then G=A:BG=A{:}B, where $A$ is abelian, $B$ is nilpotent, and $A\cap{\bf Z}(G)=1$.
@@
The proof proceeds by induction on $n$, the number of prime divisors of $|G|$. For $n=1$, the lemma holds with $A=1$, $B=G$. Suppose $n\geqslant 2$, and assume inductively that the lemma holds for groups of order with less prime divisors.
@@
with $C$ abelian, $D$ nilpotent, and $C\cap{\bf Z}(C{:}D)=1$, so that, $D$ does not centralize any Sylow subgroup of $C$.
If $G=G_{p}\times G_{p^{\prime}}$, then
$G=G_{p}{\times}(C{:}D)=(G_{p}{\times}C){:}D=C{:}(G_{p}\times D),$
which gives an expected factorization of $G$, with $A=C$, $B=G_{p}\times D$.